OpenStudy (anonymous):

How do I find the foci of this equation? x^2 − 5y^2 = 6

OpenStudy (anonymous):

divide your equation all through by 6 \[x ^{2} /6 -5/6y ^{2} = 1\] \[6^{2} - \left( 5/6 \right)^{2} = 35.3\] \[\sqrt{35.3} =5.9\] foci: (0,5.9) (0,-5.9)

OpenStudy (anonymous):

thax :)

OpenStudy (anonymous):

shouldn't the foci be (5.9, 0)?